OpenStudy (anonymous):

how many atoms are in 187 g of calcium? can someone help me pls

OpenStudy (aaronq):

use the formulas i showed you before, \(moles=\dfrac{mass}{molar\;mass}\) and \(moles=\dfrac{particles}{avogadros\;number}\)

OpenStudy (anonymous):

187/12.01 right

OpenStudy (abb0t):

Where did you get 12.01??

OpenStudy (abb0t):

That's not the molar mass for calcium

OpenStudy (anonymous):

idk i dont know how to it? can u help do it plz

OpenStudy (anonymous):

i want how many atoms are in 187 g of calcium

OpenStudy (abb0t):

Follow what the person above me first told you. Divide the mass given by the \(\sf \color{red}{molar~mass}\) of calcium. Then by avogadro's number, which is: \(\sf \color{blue}{6.022 \times 10^{23} \frac{atoms}{mol}}\)

OpenStudy (abb0t):

The molar mass can be found by looking at your periodic table, where it says "Ca" for calcium.

OpenStudy (anonymous):

so thats the answer

OpenStudy (abb0t):

No, that's not the answer. You have to solve for it. I gave you the steps

OpenStudy (anonymous):

the calcium is 40.08

OpenStudy (gtxmuqsit):

@emanuelfarias sory I see your qts now wait for 3 minutes

OpenStudy (gtxmuqsit):

\[2.0809x10^{24}\]

OpenStudy (anonymous):

@GTXMUQSIT okay just need this question how many atoms are in 187 g of calcium?

OpenStudy (gtxmuqsit):

2.0809x10 to the power 24 atoms

OpenStudy (gtxmuqsit):

@emanuelfarias

OpenStudy (anonymous):

its wrong:(

OpenStudy (gtxmuqsit):

why?

OpenStudy (anonymous):

@GTXMUQSIT

OpenStudy (anonymous):

IDK

OpenStudy (gtxmuqsit):

what's the anwer then? @emanuelfarias

OpenStudy (gtxmuqsit):

answer

OpenStudy (anonymous):

i have 5 min to finish it idk is not telling me omg

OpenStudy (anonymous):

okay so we need 187g of calcium to atoms right

OpenStudy (gtxmuqsit):

there are 2 ways to write the answer first I told ya second 20.809x10 to the power 23

OpenStudy (gtxmuqsit):

yeah first divide given mass with molar mass then multiply it with avagadro's no.

OpenStudy (anonymous):

nope it didint work

OpenStudy (anonymous):

idk but with the first u say 2.0809 is not 20.809

OpenStudy (gtxmuqsit):

@emanuelfarias buddy \[2.0809*10^{24}\]

OpenStudy (gtxmuqsit):

\[20.089*10^{23}\]

OpenStudy (gtxmuqsit):

both same thing

OpenStudy (gtxmuqsit):

@emanuelfarias

OpenStudy (abb0t):

Follow my instructions. Don't give direct answers.

OpenStudy (gtxmuqsit):

20.809x10 to the power 23 and 2.809x10 to the power 24 both are same in term of scientific notation just see decimal @emanuelfarias

OpenStudy (gtxmuqsit):

@abb0t I have proved him above how to solve

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

@GTXMUQSIT Thanks alot u were alot of help u r the best:) thank u again im going off to class

OpenStudy (gtxmuqsit):

@emanuelfarias ok welcome

OpenStudy (anonymous):

@GTXMUQSIT HEllo

OpenStudy (gtxmuqsit):

@emanuelfarias how are you?

OpenStudy (anonymous):

@GTXMUQSIT HI

OpenStudy (anonymous):

good and u